JORDAN S. ELLENBERG

Work Address: Home Address: Department of Mathematics Hibben Apartments 3-L Princeton University Faculty Road Fine Hall · Washington Road Princeton, NJ 08540 Princeton, NJ 08544-1000 (609)-688-9429 (609)-258-6467

ellenber@math.princeton.edu

Employment Instructor, Princeton University (1998-)

MSRI Postdoctoral Fellow (Fall 1999)

Education Harvard University, Cambridge, MA. Ph.D. in Mathematics, June, 1998. Dissertation under the supervision of Prof. Barry Mazur: Hilbert modular forms and the Galois representations associated to Hilbert-Blumenthal abelian varieties. Harvard University, Cambridge, MA. A.B. summa cum laude in Mathematics, June, 1993.

Research Interests  Number theory, algebraic geometry.
Publications “Endomorphisms of Jacobians”, submitted

“On the modularity of Q-curves”, submitted (with C. Skinner)

“Finiteness of torsion subschemes of Hilbert-Blumenthal abelian varieties,” to appear, Jour. für die Reine und Angew. Math. “Congruence ABC implies ABC,” to appear, Indag. Math. “The combinatorics of rewritability in finite groups,” Group theory (Granville, OH, 1992), 250– 261, World Sci. Publishing, River Edge, NJ, 1993 (with G. Sherman, L. Smithline, C. Sugar, E. Wepsic)

Awards and Prizes National Science Foundation Graduate Fellow (1994–1997) Department of Mathematics, Harvard University. Barry M. Goldwater Scholarship (1991–92) Received national scholarship for undergraduate study of mathematics. William Lowell Putnam Competition (1989–92) Placed in top six on national competition in 1990 and 1992, top ten in 1989 and 1991. U.S.A. Mathematical Olympiad (1989) Received first place. International Mathematical Olympiad (1987–89) One of six members of U.S. team; received gold medal in 1987 and 1989, silver medal in 1988.

Invited Talks Ohio State University Number Theory Seminar, April 2000

Midwest Algebraic Number Theory Day, April 2000

University of Illinois Number Theory Seminar, March 2000

Columbia University Algebraic Geometry Seminar, February 2000

MSRI Workshop: “Modularity of elliptic curves and beyond,” December 1999

Berkeley Number Theory Seminar, November 1999

Park City Math Institute Research Seminar, July 1999

Institute for Defense Analysis Colloquium, May 1999

M.I.T. Algebraic Geometry Seminar, May 1999

AMS Special Session on Galois Representations, March 1999 University of Pennsylvania Algebra Seminar, January 1999

Five College Number Theory Seminar, at University of Massachusetts, December 1998

Stanford University Number Theory Seminar, November 1998

Columbia University Number Theory Seminar, October 1998

Teaching Experience Instructor, Princeton University (1998-) Taught Math 104, Calculus, and Math 204, Advanced Linear Algebra, to Princeton undergraduates. Expository Article “The idea of a moduli space,” article for high school students, published in Imagine, May/June 1998, reprinted in Math Horizons, November 1998. Teaching Fellow (Spring 1997, Spring 1998) Taught a section of Math 21b, Linear Algebra and Differential Equations, to Harvard undergraduates. (This course meets only in section, so I had sole teaching responsibility.) Non-resident Tutor in Mathematics, Eliot House (1994–1998) Served as academic advisor for undergraduate math majors living in Eliot House at Harvard. Course Assistant (Fall 1990) Course assistant for Math 1b, Calculus, Series, and Differential Equations, at Harvard.